psirockin2:

Could someone check this math problem for me? I feel like it's wrong and I have no way to check it (only the odds are in the back of the book)

psirockin2:

https://cdn.discordapp.com/attachments/603011258970800151/688792819917127701/20200315_125422.jpg Sorry, mobile doesn't like to cooperate

satellite73:

first part is ok

satellite73:

second part not sure what you are doing. if you want to rationalize the denominator then \[\frac{1+\sqrt3}{1-\sqrt3}\]

satellite73:

oops, gone dang

AyellowBlanket:

sorry man, the page is blocked, I can't view it

AyellowBlanket:

@darkknight

InsatiableSuffering:

I can confirm that tan(7pi/12) matches with what you wrote down.

jhonyy9:

\[(1+\sqrt{3})/(1 -\sqrt{3}) = \]

jhonyy9:

multiplie the numerator and denominator by conjugate of denominator do you know it ?

AyellowBlanket:

why is the picture sideways?

jhonyy9:

how you ve wrote on your posted paper not is correct bc. \[1-\sqrt{3}\] not is the conjugate of \[1-\sqrt{3}\]

Kavithab:

It is correct